Light travels in the positive x-direction. At a particular location and time the electric field points in the positive y-direction. In which direction is the magnetic field?

a) positive z-direction

b) negative z-direction
Answer #1

As we know that the direction of light travel = direction of electric field \times direction of magnetic field ('\times' denoted cross-product)

Here,

direction of light travel = positive x direction = + i

direction of electric field = positive y direction = + j

Therefore,

+i = +j \times direction of magnetic field

==> direction of magnetic field = + k = positive z-direction ==> ANSWER-(a)